### v1

# An upper bound of the silhouette validation metric for clustering

Abstract
The silhouette coefficient summarizes, per observation, cohesion versus separation in [1,1][-1,1]; the average silhouette width ($\mathrm{ASW}$) is a common internal measure of clustering quality where higher values indicate more coveted results. However, the dataset-specific maximum of $\mathrm{ASW}$ is typically unknown, and the standard upper limit $1$ is often unattainable. In this work, we derive for each data point in a given dataset a sharp upper bound on its silhouette width. By aggregating these individual bounds, we present a canonical data-dependent upper bound on $\mathrm{ASW}$ that often assumes values well below $1$. The presented bounds can indicate whether individual data points can ever be well placed, enable early stopping of silhouette-based optimization loops, and help answer a key question: How close is my clustering result to the best possible outcome on this specific data? Across synthetic and real datasets, the bounds are provably near-tight in many cases and offer significant enrichment of cluster quality evaluation.
keywords:
Data mining , cluster analysis , internal evaluation metric , silhouette score , average silhouette width , upper bound
[…]
1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [1]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [21]. Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [42].
Among cluster validation indices, the silhouette—also known as the silhouette coefficient, silhouette score, or silhouette index—is one of the most widely used measures for assessing how well each point fits within its assigned cluster relative to the nearest neighboring cluster [74]. The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$. $\mathrm{ASW}$ values close to 1 signal well-separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned. Extensive work has examined and refined $\mathrm{ASW}$. [11] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, [67] proposed alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets. Intrinsic data characteristics can impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$. Directly comparing $\mathrm{ASW}$ values between such datasets can therefore lead to misguided conclusions.
To address this issue, we derive a data-dependent upper bound on the $\mathrm{ASW}$. This bound supplies essential context: if analysis shows that no partition can exceed, say, an $\mathrm{ASW}$ of 0.30, an empirical result of 0.29 should be regarded as essentially optimal. Such insight prevents wasted effort in pursuit of negligible gains and enables more informed evaluation of clustering quality.
A further obstacle is the combinatorial explosion of the partition space. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion, including the silhouette index, computationally prohibitive. [11] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a principled and data-aware benchmark, namely a theoretical bound, against which the quality of such locally optimal solutions can be evaluated.
Figure 1: Illustration of the silhouette coefficient calculation for $x_{1}$, showing intra-cluster distances within cluster $C_{1}$ and inter-cluster distances to cluster $C_{2}$.
In general, the major contributions of this paper are as follows:
    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$ to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass. We also derive for each data point an upper bound on its silhouette width. These per-point bounds have the potential to highlight observations that cannot be well-placed in any clustering. The bounds are computed by an $O(n^{2}\!\log n)$ algorithm that is practical for medium– to large-scale datasets. In some scenarios, when combined with empirical results, the proposed $\mathrm{ASW}$ upper bound can serve as proof that a silhouette-optimal solution has actually been found.
    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures and clustering baselines. The bound is tight or near-tight in many cases, diagnoses weak clusterability when no inherent group structure exists, tightens under minimum cluster-size constraints, and enables early stopping in silhouette-based model selection.
    All datasets, preprocessing scripts, bound-computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
The rest of this paper is organized as follows. Section 2 reviews related works. Section 3 introduces preliminaries. Section 4 details the proposed method. Section 5 presents the experimental results. Finally, section 6 summarizes the findings and outlines directions for future research.
2 Related work
Silhouette coefficientEnergy & Power Systems[62][73][93][23][27][91][51][94][40][2][19]Transportation & Mobility[60][86][8][79][37][66][58][56][4]Biomedical & Health[43][78][33][34][9][47][75][44][87][12]Food & Agriculture[25][81][32]Social & Economic[63][72][36][90][5]Materials & Chemistry[69][13][76][6][17]Environment & Earth Sciences[61][38][89][41][10][14][35][39][92][18]Sensor Networks & IoT[49][50][64][65][85]Big Data[70][83][82][16]Cluster Validity & Selection[45][46][11][55][15][30][7]Algorithm Development[48][80][59][71][90][8][79][49][50][22][29][31][28][26][32]Feature & Dim-Reduction[75][24][95][16]
Figure 2: Taxonomy of how the silhouette coefficient is employed across research fields.
The silhouette coefficient [74] is a distance-based internal validity index that quantifies how well each object fits its assigned cluster relative to its nearest alternative. Fig. 1 illustrates the construction. For a point $x_{i}\in C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$ (its average dissimilarity to members of $C_{\ell}$) and nearest-cluster separation by $b(x_{i})$ (its average dissimilarity to the closest rival cluster); see Sec.  3 for formal definitions. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own.
Aggregating over points gives the average silhouette width (ASW), a scale-free summary of partition quality (Sec.  3). In practice, ASW is often used for internal model selection by choosing the number of clusters $K$ that maximizes it; rule-of-thumb interpretations treat larger ASW as evidence of well-separated, compact clusters, and values near zero or negative as signs of overlap or a poor choice of $K$. Despite numerous extensions—e.g., to categorical data [20], fuzzy clustering [53], hierarchical clustering [63], model-based clustering [35], spectral clustering [87], spherical clustering [18], and optimization-based methods [88]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
3 Preliminaries
Table 1: A toy dataset of five samples with two features.
Point $Att_{1}$ $Att_{2}$
$x_{1}$ 1.0 2.0
$x_{2}$ 2.0 1.0
$x_{3}$ 1.5 2.5
$x_{4}$ 6.0 2.0
$x_{5}$ 6.0 3.0
Figure 3: Scatter plot of the toy dataset.
We consider a dataset of $n$ observations described by $m$ features. Throughout, we fix an integer $n\geq 4$ and index the observations by the set $\{1,\dots,n\}$. For example, Tab. 3 and Fig. 3 present a dataset consisting of five data points, each described by two features. We use this toy dataset for illustrative examples in this section.
When applicable, the samples $\{\mathbf{x}_{i}\}_{i=1}^{n}$ of a dataset can be stored in a matrix of the form
$X\;=\;[x_{ij}]_{1\leq i\leq n,\;1\leq j\leq m}\;\in\;\mathbb{R}^{n\times m},\qquad\mathbf{x}_{i}:=(x_{i1},\dots,x_{im})^{\top}\in\mathbb{R}^{m}.$
We deal with datasets where there is a measure of dissimilarity (or distance) between each pair of data points. All pairwise dissimilarities are contained in the datasets dissimilarity matrix.
Definition 1 (Dissimilarity matrix).
A matrix $\Delta\in\mathbb{R}^{n\times n}$ is a dissimilarity matrix if
  1. (a)
    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non-negative),
  2. (b)
    $\Delta_{ii}=0$ for all $i$ ($\Delta$ has zero diagonal), and
  3. (c)
    $\Delta_{ij}=\Delta_{ji}$ for all $i,j$ ($\Delta$ is symmetric).
The triangle inequality is not required. We additionally exclude matrices containing a row of all zeros.
Figure 4: The first 27 Bell numbers, which enumerate the partitions of an $n$-element set. This rapid growth highlights the combinatorial explosion of clustering possibilities as $n$ increases.
For instance, if the Euclidean distance is used, then $\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2}$. The dissimilarity matrix corresponding to the dataset shown in Fig. 3 is:
$\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2},\qquad\Delta=\begin{pmatrix}0&1.414&0.707&5.000&5.099\\ 1.414&0&1.581&4.123&4.472\\ 0.707&1.581&0&4.528&4.528\\ 5.000&4.123&4.528&0&1.000\\ 5.099&4.472&4.528&1.000&0\end{pmatrix}$
We remark that the upper bound proposed in Sec. 4 depends only on the dissimilarity matrix of the dataset; no observation vectors are required. Hence, it is just as data-agnostic as the $\mathrm{ASW}$ itself.
For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order. We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$). In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$. Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
The sorted off-diagonal dissimilarity matrix for the toy dataset is shown in Tab. 2.
Table 2: Row-wise sorted off-diagonal distances $\hat{\Delta}$ corresponding to the toy dataset.
Point $\hat{\Delta}_{i1}$ $\hat{\Delta}_{i2}$ $\hat{\Delta}_{i3}$ $\hat{\Delta}_{i4}$
$x_{1}$ 0.707 1.414 5.000 5.099
$x_{2}$ 1.414 1.581 4.123 4.472
$x_{3}$ 0.707 1.581 4.528 4.528
$x_{4}$ 1.000 4.123 4.528 5.000
$x_{5}$ 1.000 4.472 4.528 5.099
Now, let us formally define the concept of a clustering of the dataset $\{1,\ldots,n\}$.
Definition 2 (Clustering).
Let $K\in\{2,\ldots,n\}$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subsetneq\{i\}_{i=1}^{n}$ for any $I$, and
[…]
The total number of clusterings of an $n$-element set grows according to the Bell numbers (Fig. 4), illustrating the combinatorial complexity of the clustering problem.
A popular clustering method is the K-means algorithm, which we will use in Sec. 5. Its objective is to minimize the within-cluster sum of squares (or variance).
Definition 3 ($K$-means [57]).
When the desired number of clusters $K$ is specified, the classical $\mathbf{K}$-means problem seeks
$\operatorname*{argmin}_{\mathcal{C}_{K}}\sum_{I=1}^{K}\;\sum_{i\in C_{I}}\bigl{\lVert}\mathbf{x}_{i}-\boldsymbol{\mu}_{I}\bigr{\rVert}_{2}^{2},\qquad\boldsymbol{\mu}_{I}\coloneqq\frac{1}{|C_{I}|}\sum_{i\in C_{I}}\mathbf{x}_{i}.$
Figure 5: Clustering results using $K$-means on the dataset in Fig. 3.
For instance, Fig. 5 illustrates the result of applying $K$-means clustering to the toy dataset with $K=2$ clusters.
Alternative objectives such as linkage-based agglomerative clustering may be formulated directly in terms of $\Delta$, but the underlying aim remains to form internally cohesive and externally separated groups. For instance, Fig.6 shows a dendrogram obtained from performing hierarchical clustering on the dissimilarity matrix corresponding to the toy dataset.
Figure 6: Hierarchical clustering on the distance matrix of data points in Fig. 3.
We will now define the silhouette, which was introduced in 1987 by P.J. Rousseeuw [74]. We employ similar notation.
Definition 4 (Silhouette).
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is 0 if $|C_{I}|=1$; otherwise it is
$s(i|\mathcal{C}_{K},\Delta)\coloneqq\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$
where
$a(i)\coloneqq\frac{1}{|C_{I}|-1}\sum_{j\in C_{I}\setminus\{i\}}\Delta_{ij},\;\;\;\text{and}\;\;\;b(i)\coloneqq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right).$
The average silhouette width (ASW) is defined as
$ASW(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i|\mathcal{C}_{K},\Delta).$
Note that the silhouette width of $i$ may be written as
$s(i|\mathcal{C}_{K},\Delta)=\begin{cases}1-a(i)/b(i)\text{ if }a(i)<b(i),\\ 0\text{ if }a(i)=b(i),\\ b(i)/a(i)-1\text{ if }a(i)>b(i).\end{cases}$
Consequently, $1\leq s(i\,|\,\mathcal{C}_{K},\Delta)\leq 1$.
To illustrate with an example, the silhouette value of each data point from the clustering results shown in Fig. 5 is displayed in Tab. 3 and Fig. 7. The corresponding $\operatorname{ASW}$ is $0.7512$, indicating well-separated and internally cohesive clusters.
Table 3: Silhouette components for each point in the toy dataset after applying K-means clustering with $K=2$.
$i$ $a(i)$ $b(i)$ $s(i)$
$x_{1}$ 1.061 5.050 0.790
$x_{2}$ 1.498 4.298 0.652
$x_{3}$ 1.144 4.528 0.747
$x_{4}$ 1.000 4.550 0.780
$x_{5}$ 1.000 4.700 0.787
Figure 7: Silhouette plots of clustering results from Fig. 5.
Individual silhouette widths provide a valuable graphical diagnostic (see Fig. 7), but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [54]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, the limitation often stems from the intrinsic structure of the dissimilarity data rather than a suboptimal choice of algorithm. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$? To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
Definition 5 (Silhouette-optimal clustering).
Let $\Delta$ be the dissimilarity matrix of the dataset $\{1,\ldots,n\}$. We define
$\mathfrak{C}\coloneqq\{\mathcal{C}_{K}:\mathcal{C}_{K}\text{ is a $K$-clustering of $\{1,\ldots,n\}$}\}$
to be the set of all possible clusterings. Any
$\mathcal{C}^{*}\in\operatorname*{argmax}_{\mathcal{C}_{K}\in\mathfrak{C}}\{\mathrm{ASW}(\mathcal{C}_{K},\Delta)\}$
is called a silhouette-optimal clustering.
Once again, recall the toy data (Tab. 3) and its dissimilarity matrix $\Delta$. We evaluate $\operatorname{ASW}(\mathcal{C}_{K},\Delta)$ for every admissible clustering with $K\in\{2,3,4,5\}$ (all set partitions except the trivial K=1K{=}1 case). Among the 51 candidates  [^3] The fifth Bell number $B_{5}=52$ and excluding K=1K{=}1 gives $51$., the maximizer is unique:
$\mathcal{C}^{\ast}\;=\;\bigl{\{}\{x_{1},x_{2},x_{3}\},\;\{x_{4},x_{5}\}\bigr{\}}\quad\text{with}\quad\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)\;=\;0.7512,$
which matches the per-point silhouettes reported in Tab. 3 and Fig. 7. Finally, the best achievable $\operatorname{ASW}$ within each fixed $K$ is depicted in Tab. 4.
Table 4: Optimal $\operatorname{ASW}$ within each $K$ on the toy dataset.
$K$ Best partition Best ASW
2 $\{x_{1},x_{2},x_{3}\}\mid\{x_{4},x_{5}\}$ 0.7512
3 $\{x_{1},x_{3}\}\mid\{x_{2}\}\mid\{x_{4},x_{5}\}$ 0.5173
4 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4},x_{5}\}$ 0.3068
5 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4}\}\mid\{x_{5}\}$ 0.0000
4 The proposed upper bound of the silhouette
4.1 Overview
We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed. For each data point, we derive a sharp upper bound of its silhouette width. Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and moreover, there exists a clustering in which the bound is exactly achieved. Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and supports optional minimum-cluster-size constraints to make it even sharper.
4.2 Derivation
Consider an arbitrary clustering $\mathcal{C}_{K}$ of the dataset $\{1,\ldots,n\}$ with dissimilarites $\Delta$. We will now formalize the upper bound of the silhouette width for point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)$.
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the other data points in the same cluster is not smaller than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning via the $\Lambda$-quotient of point $i$.
Definition 6.
Let $\Lambda\in\mathbb{N}$ with $1\leq\Lambda\leq n-1$. The $\Lambda$-quotient of data point $i$ is
$q(i,\Delta,\Lambda)\coloneqq\begin{cases}\dfrac{n-\Lambda}{\Lambda-1}\cdot\dfrac{\sum_{j=1}^{\Lambda-1}\hat{\Delta}_{ij}}{\sum_{j=\Lambda}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }\Lambda>1,\\[4.30554pt] 1,&\textnormal{if }\Lambda=1.\end{cases}$
Recall the following basic fact.
Lemma 1.
Let $A=\{a_{1},\ldots,a_{N}\}$ be a collection of $N$ real values, and consider a partition of $A$ into subsets $A_{1},\ldots,A_{M}$. Then
$\min_{L}\frac{1}{|A_{L}|}\sum_{a\in A_{L}}a\leq\frac{1}{|A|}\sum_{a\in A}a.$
Proof.
Let $\mu=\frac{1}{|A|}\sum_{a\in A}a$ and $\mu_{L}=\frac{1}{|A_{L}|}\sum_{a\in A_{L}}a$. We have $\mu=\sum_{L=1}^{M}\frac{|A_{L}|}{|A|}\mu_{L}$, and since $\sum|A_{L}|=|A|$, the statement holds. ∎
Proposition 1.
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{i\}_{i=1}^{n}$ with dissimilarites $\Delta$. Let $i$ represent any data point and let $C_{I}\in\mathcal{C}_{K}$ denote the cluster that contains $i$. Furthermore, let
$f(i,\Delta)\coloneqq\min_{1\leq\Lambda\leq n-1}\{q(i,\Delta,\Lambda)\}$
denote a minimal $\Lambda$-quotient. Then
$s(i|\mathcal{C}_{K},\Delta)\leq 1-f(i,\Delta).$
Proof.
Suppose $q(i,\Delta,2)>1$. Then $(n-2)\hat{\Delta}_{i1}>\sum_{j=2}^{n-1}\hat{\Delta}_{ij}$, which contradicts $\hat{\Delta}_{i1}\leq\hat{\Delta}_{ij}$ for $j\geq 2$. Hence, $0\leq f(i,\Delta)\leq 1$. This proves the statement if $|C_{I}|=1$ and for cases $a(i)\geq b(i)$ when $|C_{I}|>1$. Consider $q(i,\Delta,|C_{I}|)$. It is clear that $\sum_{j=1}^{|C_{I}|-1}\hat{\Delta}_{ij}/(|C_{I}|-1)\leq a(i)$. It remains to argue that $\sum_{j=|C_{I}|}^{n-1}\hat{\Delta}_{ij}/(n-|C_{I}|)\geq b(i)$. By Lemma 1, we have
$\frac{1}{n-|C_{I}|}\sum_{j=|C_{I}|}^{n-1}\hat{\Delta}_{ij}$ $\geq\frac{1}{n-|C_{I}|}\sum_{j\in\bigcup_{J\neq I}C_{J}}\Delta_{ij}$
$\geq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right)=b(i).$
Thus, $f(i,\Delta)\leq q(i,\Delta,|C_{I}|)\leq a(i)/b(i)$ and $1-f(i,\Delta)\geq 1-a(i)/b(i)$. ∎
Taking the mean of all silhouette width upper bounds from Proposition 1 results in the following inequality, which is the main observation of this paper:
$ASW(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
We remark that the right-hand side in 1 depends only on the underlying data $\Delta$. Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$. Moreover, since any $\Lambda$-quotient naturally corresponds to a $2$-clustering of the dataset, the bound in Proposition 1 is indeed sharp.
Finally, imposing the narrower definition
$f_{\kappa}(i,\Delta)\coloneqq\min_{\kappa\leq\Lambda\leq n-\kappa}\{q(i,\Delta,\Lambda)\}$
for some $\kappa\in\{2,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$. Often, as we shall see, this leads to strictly tighter limits.
4.3 Algorithm
Input: $\Delta\in\mathbb{R}^{n\times n}$;
  // Dissimilarity matrix
1
Input: $\kappa$;
  // Smallest allowed cluster size, $1\leq\kappa\leq\lfloor n/2\rfloor$
2
Output: $A\in\mathbb{R}^{n}$;
  // Array of individual upper bounds
3
4 Initialize array $A\leftarrow[\;]$ ;
5 Compute $\hat{\Delta}$ from $\Delta$;
6 for $i=1,\ldots,n$ do
7 $y\leftarrow\sum_{j=\kappa}^{n-1}\hat{\Delta}_{i,j}$ ;
8 $x\leftarrow 0$ ;
 $q\leftarrow 1$ ;
    // Initialize $\Lambda$-quotient
9 if $\kappa>1$ then
10    $x\leftarrow\sum_{j=1}^{\kappa-1}\hat{\Delta}_{i,j}$ ;
11    $q\leftarrow\frac{x}{\kappa-1}\big{/}\frac{y}{n-\kappa}$;
12
13 for $\Lambda=\kappa+1,\ldots,n-\kappa$ do
14    $x\leftarrow x+\hat{\Delta}_{i,\Lambda-1}$;
15    $y\leftarrow y-\hat{\Delta}_{i,\Lambda-1}$;
16    $q_{\text{candidate}}\leftarrow\dfrac{x}{\Lambda-1}\big{/}\dfrac{y}{n-\Lambda}$;
17    if $q_{\text{candidate}}<q$ then
       $q\leftarrow q_{\text{cand}}$ ;
          // Update minimal $\Lambda$-quotient
18
19
20  Append 1q1-q to AA;
21
return AA
Algorithm 1 Silhouette upper bound (pointwise)
The workflow of the proposed upper bound on the ASW is shown in Fig.8. Following this workflow, Algo. 1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the (restricted) solution space $\min_{I}|C_{I}|\geq\kappa$. The steps are as follows: (1) For each row $i$ of $\Delta$, we sort the off-diagonal entries increasingly to obtain $\hat{\Delta}_{i,1},\hat{\Delta}_{i,2},\ldots$, the nearest-to-farthest distances from $i$. (2) For $i$, we initialize the running sums $x$ and $y$ so that $x/(\kappa-1)$ is its mean distance to its $\kappa-1$ nearest points, and $y/(n-\kappa)$ is its mean distance to its $n-\kappa$ farthest points. (3) Starting from $\Lambda=\kappa+1$, we update these running sums as $\Lambda$ increases, thereby scanning all admissible values $\Lambda\in\{\kappa,\ldots,n-\kappa\}$. For each $\Lambda$, we form the $\Lambda$-quotient $q(i,\Delta,\Lambda)$ and keep the minimum. (4) Finally, we record 1q1-q as the upper bound for $s(i\,|\,\mathcal{C}_{K},\Delta)$. The array $A$ collects these values across all points, and its average equals the dataset-level upper bound in Eq. (1) when $\kappa=1$.
Complexity
Forming $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\mathrm{ASW}$; $\mathcal{O}(n^{2})$ [54].
Output summaries
Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo. 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). To reflect the spread of the individual bounds, we denote by $\texttt{minUB}(\Delta)$ and $\texttt{maxUB}(\Delta)$ the minimum and maximum values of $A$, respectively. Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects. Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space.
Remark
The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$; its tightness is data-dependent. Nonetheless, it provides an enlightening piece of information that merits inclusion in analyses.
4.4 Example
Recall the toy data from Tab. 3. We have $n=5$, so $\Lambda\in\{1,2,3,4\}$. The values $q(i,\Delta,\Lambda)$ for all instances are shown in Tab. 5. Consider the K=2K{=}2 clustering from Fig. 5: $C_{1}=\{x_{1},x_{2},x_{3}\}$ and $C_{2}=\{x_{4},x_{5}\}$. Tab. 6 compares each silhouette width with the corresponding upper bound, illustrating that $s(i)\leq 1-f(i,\Delta)$ for all $i$, with equality for $x_{2},x_{4},x_{5}$ and a small strict headroom for $x_{1}$ ($0.026$) and $x_{3}$ ($0.054$). Moreover, we have
$\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)=0.7512\;\leq\;0.7672=1-\frac{1}{n}\sum_{i=1}^{n}f(i,\Delta).$
Table 5: All possible values of $q(i,\Delta,\Lambda)$ for each point $i$ in the toy dataset.
Point $q(\cdot,1)$ $q(\cdot,2)$ $q(\cdot,3)$ $q(\cdot,4)$
$x_{1}$ 1.000 0.184 0.210 0.466
$x_{2}$ 1.000 0.417 0.348 0.531
$x_{3}$ 1.000 0.199 0.253 0.502
$x_{4}$ 1.000 0.220 0.538 0.643
$x_{5}$ 1.000 0.213 0.568 0.654
Table 6: Showing $s(i)\leq 1-f(i,\Delta)$ for each point in the toy dataset.
$i$ $|C_{I}|$ $\Lambda^{\star}=\arg\min_{\Lambda}q(i,\Delta,\Lambda)$ $f(i,\Delta)$ $1-f(i,\Delta)$ $s(i)$
$x_{1}$ 3 2 0.184 0.816 0.790
$x_{2}$ 3 3 0.348 0.652 0.652
$x_{3}$ 3 2 0.199 0.801 0.747
$x_{4}$ 2 2 0.220 0.780 0.780
$x_{5}$ 2 2 0.213 0.787 0.787
Figure 8: Workflow of the proposed upper bound on the $\mathrm{ASW}$.
4.5 Enabling early stopping of optimization loops through the upper bound
We now demonstrate how the data–dependent upper bound introduced in Eq. (1) can be used to accelerate standard ASW based optimization loops. In many practical applications, selecting the number of clusters involves iterating over $K$ and evaluating the output of the clustering algorithm $\mathcal{A}(K)$ in order to maximize the ASW. When the dataset is large, each iteration can be computationally demanding, and exhaustive searches over all candidate $K$ quickly become demanding. By incorporating the upper bound into this procedure, we can enable early stopping: if the current ASW is provably within a user–specified relative tolerance of the global optimum, the loop halts, saving potentially substantial computation while still guaranteeing near–optimal clustering performance within the search space. This type of procedure is described in more detail in Algo. 2.
Input: $\Delta\in\mathbb{R}^{n\times n}$;
  // Dissimilarity matrix
1
Input: $\epsilon$;
  // Relative error tolerance, $0\leq\epsilon\leq 1$
2
Input: $\tau$;
  // Clusterability threshold, $0\leq\tau\leq 1$
3
Input: $\mathcal{A}(K),K_{\max}$;
  // Clustering algorithm, max. no. clusters
4
Output: ASW-optimal clustering of $\Delta$ (within search space)
Compute $\mathrm{UB}(\Delta)$;
  // Apply Algo. 1
5 if $\mathrm{UB}(\Delta)\leq\tau$ then // Abort if dataset is insufficiently clusterable
6 return NULL
7optimal clustering $\leftarrow$ NULL ;
8 optimal ASW $\leftarrow$ -1 ;
9 for $K=2,\ldots,K_{\max}$ do
[…]
11      optimal clustering $\leftarrow\mathcal{A}(K)$ ;
12      optimal ASW $\leftarrow\mathrm{ASW}(\mathcal{A}(K),\Delta)$ ;
13    if $\frac{\mathrm{UB}(\Delta)-\text{optimal ASW}}{\mathrm{UB}(\Delta)}<\epsilon$ then
14       return optimal clustering
15
16
return optimal clustering
Algorithm 2 Early stopping in ASW-based optimization
Workflow
The procedure integrates the proposed upper bound into a standard ASW optimization loop. At each step, the current best clustering is compared to the global upper bound. If the relative gap falls below the user-specified tolerance $\epsilon$, the algorithm terminates early, certifying that no further search can improve ASW by more than $\epsilon\%$ of the true global optimum. In the worst case, the condition is never met, and the loop reverts to the standard exhaustive search over $K=2,\ldots,K_{\max}$, ensuring no loss in clustering quality.
Remark
If minimal cluster size constraints are of interest, the procedure can be modified by replacing the initial call to $\mathrm{UB}(\Delta)$ with the restricted version $\mathrm{UB}_{\kappa}(\Delta)$. This enables early stopping under user-specified requirements. We also emphasise that clusterability in this context is interpreted solely through the lens of the silhouette coefficient. While this provides a useful and practical criterion for our purposes, it is by no means a universal or definitive notion of clusterability—many alternative perspectives exist in the literature.
Recall the toy dataset summarized in Tab. 4. In Sec.  4.4, we established that the corresponding upper bound is $\texttt{UB}(\Delta)=0.7672$. Let $\widehat{S}_{K}$ denote the empirically observed $\mathrm{ASW}$ value for $K$ clusters. For the toy dataset, these values are:
$\widehat{S}_{2}=0.7512,\quad\widehat{S}_{3}=0.5173,\quad\widehat{S}_{4}=0.3068,\quad\widehat{S}_{5}=0.0000.$
The performance metric of interest is the relative error, defined as
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)},$
where $C^{*}$ denotes the clustering achieving the global maximum ASW. Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}_{K}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical $\widehat{S}_{K}$ is guaranteed to be within $\epsilon$ of the true maximum ASW.
Table 7: Early stopping on the toy dataset.
$K$ $\widehat{S}_{K}$ $\texttt{UB}(\Delta)$ worst-case relative error
[…]
As illustrated by Tab. 7, applying Algo. 2 with relative error tolerance $\epsilon=0.05$ for the toy dataset would halt the process already at $K=2$, thereby avoiding unnecessary evaluations for $K=3,4,5$. In this way, the upper bound allows substantial savings while still certifying near-optimal performance.
5 Experiments, results & discussion
5.1 Experimental datasets
Table 8: Datasets used in our experiments and their characteristics. The Metric(s) column indicates the distance metrics used in the experiments.
ID Dataset Name Instances Features Classes Field Metric(s)
1 gene Gene expressions 801 20531 Unknown Genomics Euclidean, correlation
2 ceramic Ceramic samples 88 17 Unknown Archaeology Euclidean
3 relig. texts Religious texts 589 8266 Unknown Humanities Cosine, Euclidean, Jaccard
4 conf. papers Conference papers 5804 11463 Unknown Computer science Cosine, Euclidean, Jaccard
5 wdbc Breast Cancer Wisconsin Diagnostic 569 30 2 Healthcare Euclidean
6 wisc Breast Cancer Wisconsin Original 699 9 2 Healthcare Euclidean
7 ecoli Ecoli 336 7 8 Biology Euclidean
8 wine Wine 178 13 3 Chemistry Euclidean
9 iono Ionosphere 351 34 2 Physics Euclidean
10 zoo Zoo 101 16 7 Biology Euclidean
11 iris Iris 150 4 3 Botany Euclidean
12 segment Image Segmentation 2310 19 7 Computer science Euclidean
13 dermatology Dermatology 366 34 6 Healthcare Euclidean
14 balance-scale Balance Scale 625 4 3 Psychology Euclidean
15 heart-statlog Statlog (Heart) 270 13 2 Healthcare Euclidean
16 thy New Thyroid 215 5 3 Healthcare Euclidean
We evaluate the upper bound by applying it to a variety of datasets, both synthetic and real-life. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. (For some parts of the experiments we use the Jaccard distance, which means the data is treated in binary form and similarity is based on the amount of shared attributes.)
An essential part of the evaluation is to establish a tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$. To this end, we adopt an empirical approach: we apply several clustering algorithms to each dataset and take the largest observed $\mathrm{ASW}$ as the lower bound. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
Figure 9: Average silhouette width $\mathrm{ASW}$ vs. number of clusters ($K$). Results on the synthetic dataset 400-64-5-6 for $K$-means, PAM (PAMSIL), and Hierarchical clustering with weighted and single linkage. For each method and each chosen $K$, we compute the resulting $\mathrm{ASW}$. The black dashed line marks the theoretical upper bound $\mathrm{UB}(\Delta)=0.376$; the red dashed line marks the upper bound when a minimum cluster size of $80$ is enforced ($0.282$). The maximum observed $\mathrm{ASW}$ occurs at K=5K=5—a partition into five equal-sized clusters of 80 elements—and is achieved by $K$-means, Hierarchical (weighted linkage), and PAMSIL.
Synthetic datasets
We follow [3] in notation and approach. Synthetic datasets are generated using the make_blobs function from scikit-learn [68]. Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion.
Real-life datasets
We use sixteen datasets spanning unlabeled (IDs 1–4) and labeled (IDs 5–16) settings (Tab. 8). The unlabeled datasets are named gene expressions, ceramic samples, religious texts, and conference papers. For these, we search over $K$ for different algorithms to maximize $\operatorname{ASW}$. The labeled datasets consist of twelve standard benchmarks across multiple domains-wdbc, wisc, ecoli, wine, iono, zoo, iris, segment, dermatology, balance-scale, heart-statlog, thy-ranging from 150 to 2,310 instances, 4 to 34 features, and 2 to 8 classes. For these datasets, we use the ground-truth class count to set $K$ (e.g., $K=8$ for ecoli), run $K$-medoids to obtain a $K$-clustering, and compare its $\operatorname{ASW}$ with the bound; we also repeat a search over varying $K$ to maximize $\operatorname{ASW}$, mirroring the unlabeled protocol.
Table 9: Experimental results on synthetic data. Each dataset is denoted as n_samples-n_features-centers-cluster_std. For Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid), we report the average silhouette width (ASW); the selected number of clusters $K$ is shown in parentheses. For $K$-medoids, $\dagger$ indicates that PAMSIL was used; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime. The last three columns summarize the output of Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively). All results use Euclidean distance.
Dataset Metric
Weighted
$K$-means
$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
400-64-5-6 Euclidean $0.249$ (5) $0.209$ (4) $0.249$ (5) $0.249^{\dagger}$ (5) $0.376$ $0.276$ $0.467$
400-64-2-2 Euclidean $0.673$ (2) $0.673$ (2) $0.673$ (2) $0.673^{\dagger}$ (2) $0.673$ $0.619$ $0.718$
400-128-7-3 Euclidean $0.522$ (7) $0.522$ (7) $0.522$ (7) $0.522^{\dagger}$ (7) $0.566$ $0.514$ $0.619$
1000-161-2-13 Euclidean $0.084$ (2) $0.068$ (2) $0.084$ (2) $0.046$ (3) $0.182$ $0.137$ $0.256$
1000-300-5-2 Euclidean $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.678$ $0.651$ $0.706$
10000-32-20-2 Euclidean $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.774$ $0.681$ $0.847$
10000-1024-20-4 Euclidean $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.454$ $0.425$ $0.488$
Table 10: Experimental results on real unlabeled data. For each dataset (the distance/dissimilarity used is shown in the “Metric” column), we report the average silhouette width (ASW) for Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid); the selected number of clusters $K$ is given in parentheses. For $K$-medoids, $\dagger$ indicates PAMSIL; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime. When the metric is incompatible with $K$-means (e.g., correlation/cosine or Jaccard), the entry is marked N/A. The last three columns summarize the output of Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively).
Dataset Metric
Weighted
$K$-means
$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
Gene expressions Correlation $0.345$ (10) $0.332$ (2) N/A $0.395^{\dagger}$ (7) $0.657$ $0.325$ $0.865$
Religious texts Cosine $0.082$ (234) $0.049$ (2) N/A $0.087^{\dagger}$ (23) $0.349$ $0.079$ $0.810$
Conference papers Cosine $0.188$ (2) $0.162$ (2) N/A $0.126$ (2) $0.473$ $0.118$ $1.0$
Religious texts Euclidean $0.846$ (2) $0.846$ (2) $0.428$ (3) $0.846^{\dagger}$ (2) $0.846$ $0.106$ $0.895$
Ceramic samples Euclidean $0.547$ (2) $0.591$ (2) $0.584$ (2) $0.584^{\dagger}$ (2) $0.855$ $0.561$ $0.950$
Conference papers Euclidean $0.418$ (2) $0.418$ (2) $0.067$ (2) $0.384$ (2) $0.439$ $0.227$ $1.0$
Gene expressions Euclidean $0.207$ (2) $0.188$ (2) $0.228$ (6) $0.230^{\dagger}$ (8) $0.419$ $0.174$ $0.624$
Religious texts Jaccard $0.031$ (106) $0.017$ (569) N/A $0.024^{\dagger}$ (24) $0.119$ $0.040$ $0.719$
Conference papers Jaccard $0.160$ (2) $0.160$ (2) N/A $0.160$ (2) $0.170$ $0.038$ $1.0$
Table 11: Early stopping on real data. Early stopping results (Algo. 2) with Euclidean dissimilarity. Best ASW per row is bold; corresponding $K$ is bold.
Dataset
$K$
cand.
$\epsilon$
Best ASW
($K$-means)
Best ASW
($K$-medoids)
UB
Best $K$
($K$-means)
Best $K$
($K$-medoids)
Worst-case
relative error
($K$-means)
Worst-case
relative error
($K$-medoids)
Early stop
($K$-means)
Early stop
($K$-medoids)
gene $[2,15]$ 0.15 0.228 0.230 0.419 6 8 0.46 0.45 false false
ceramic $[2,30]$ 0.35 0.584 0.584 0.855 2 2 0.32 0.32 true true
relig. texts $[2,15]$ 0.15 0.428 0.846 0.846 3 2 0.49 0.0 false true
conf. papers $[2,15]$ 0.15 0.067 0.384 0.439 2 2 0.85 0.13 false true
wdbc $[2,15]$ 0.15 0.345 0.661 0.708 2 2 0.51 0.07 false true
wisc $[2,15]$ 0.15 0.574 0.574 0.844 2 2 0.32 0.32 false false
ecoli $[2,15]$ 0.15 0.398 0.836 0.854 4 2 0.53 0.02 false true
wine $[2,30]$ 0.35 0.285 0.295 0.626 3 5 0.55 0.53 false false
iono $[2,15]$ 0.15 0.293 0.413 0.691 6 2 0.58 0.40 false false
zoo $[2,30]$ 0.35 0.547 0.547 0.839 30 22 0.35 0.35 true true
iris $[2,30]$ 0.35 0.580 0.580 0.878 2 2 0.34 0.34 true true
segment $[2,15]$ 0.15 0.357 0.778 0.929 7 2 0.62 0.16 false false
dermatology $[2,15]$ 0.15 0.277 0.357 0.609 3 2 0.55 0.42 false false
balance-scale $[2,15]$ 0.15 0.219 0.214 0.737 8 8 0.70 0.71 false false
heart-statlog $[2,30]$ 0.35 0.169 0.322 0.593 3 2 0.72 0.46 false false
thy $[2,30]$ 0.35 0.577 0.734 0.835 2 2 0.31 0.12 true true
5.2 Benchmarking algorithms
To cluster the datasets, we employ three popular methods: $K$-means, $K$-medoids, and Hierarchical agglomerative clustering. These methods serve as baselines to contextualize the performance of the proposed upper bound.
Because our goal is to obtain the highest possible $\mathrm{ASW}$, we evaluate multiple values of $K$ and report the one that yields the largest $\mathrm{ASW}$.
$K$-means clustering
We use the implementation of $K$-means from scikit-learn. For each dataset, the algorithm requires the number of clusters $K$.
$K$-medoids clustering
We use the Python package $K$-medoids [77] and, in particular, two medoid-based algorithms: PAMSIL and PAMMEDSIL, both introduced in [52]. PAMSIL aims to directly optimize $\mathrm{ASW}$ but is computationally expensive. PAMMEDSIL is a faster proxy that optimizes a medoid-based version of the silhouette; while expedient, it generally results in a lower $\mathrm{ASW}$ than PAMSIL. To balance accuracy and run time, we use PAMSIL for datasets with $n<1000$ and switch to PAMMEDSIL for $n\geq 1000$. This threshold was chosen empirically to keep computation tractable while preserving $\mathrm{ASW}$ quality.
Hierarchical agglomerative clustering (HAC)
For hierarchical clustering, we use the SciPy implementation [84]. To measure inter-cluster distances when forming the hierarchy, we consider both single and weighted linkage.
Figure 10: Silhouette plots for the 4 real unlabeled datasets. $K$-means clusters; ASW compared against $\texttt{UB}(\Delta)$ and the size-constrained bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size.
Figure 11: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ fixed to the ground truth; ASW compared against $\texttt{UB}(\Delta)$.
Figure 12: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ chosen by unrestricted optimization; ASW compared with $\texttt{UB}(\Delta)$, illustrating the best empirical quality without fixing $K$ to ground truth.
Figure 13: Unlabeled datasets: violin plots of pointwise silhouette upper bounds versus minimum cluster size $\kappa$ (up to $\sim 10\%$ of $n$). The violins summarize distributions across data points (means, i.e. $\mathrm{ASW}$ upper bounds, in red). The upper bounds decline as $\kappa$ increases, showing that size constraints lead to sharper limits.
5.3 Results and discussion
Fig. 9 presents the $\mathrm{ASW}$ obtained by different clustering algorithms on the synthetic dataset 400-64-5-6. K-means achieves its best result with five clusters of equal size ($80$ objects each), reaching an $\mathrm{ASW}$ of $0.249$. Hierarchical clustering with weighted linkage also favors five clusters, whereas single linkage attains its maximum $\mathrm{ASW}$ ($0.209$) at four clusters. K-medoids reaches the same output as K-means and weighted linkage. The theoretical upper bound without restrictions on cluster size ($\kappa=1$) is $0.376$ (black dashed line). When enforcing a minimum cluster size of $\kappa=80$, Algo. 1 yields a reduced upper bound of $0.282$ (red dashed line). Thus, among solutions with no cluster smaller than $80$ elements, the five-cluster partition found by e.g. K-means is guaranteed to be within $12\%$ of the global optimum.
Tab. 5.1 reports the remaining results on the synthetic datasets. For each algorithm, we show the best observed $\mathrm{ASW}$, with the selected number of clusters in parentheses. The last three columns summarize the upper bounds from Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ bound), $\texttt{minUB}(\Delta)$ (smallest individual bound), and $\texttt{maxUB}(\Delta)$ (largest individual bound). For dataset 400-64-2-2, the upper bound matches the achieved $\mathrm{ASW}$ for all algorithms, certifying global optimality. Dataset 1000-161-2-13 has a significantly low upper bound ($\mathrm{UB}(\Delta)=0.182$), implying that no clustering can exceed an $\mathrm{ASW}$ of 0.182; in such cases, alternative quality measures may be more appropriate, or the data may be effectively non-clusterable.
Tab. 5.1 reports results on the real unlabeled datasets and mirrors the layout of Tab. 5.1. By default, we search over $K\in\{2,\ldots,n\}$, where $n$ is the number of samples. When necessary for reasonable run time, however, we limit the search at $K\leq 30$. Entries marked N/A indicate that $K$-means is incompatible with the metric, e.g., correlation, cosine or Jaccard. For Euclidean “Religious texts”, Euclidean “Conference papers”, and Jaccard “Conference papers”, the hierarchical methods are within $0.021$ of the upper bound $\mathrm{UB}(\Delta)$—so any further improvement in ASW over the reported values is at most $0.021$. Several datasets have small $\texttt{minUB}(\Delta)$ (e.g., Cosine “Religious texts” with $\texttt{minUB}(\Delta)=0.079$), certifying that at least one observation cannot achieve a silhouette width exceeding that value under any clustering. In general, when $\texttt{minUB}(\Delta)$ is near zero, it may be reasonable to treat the corresponding observations as noise or remove them from the analysis.
We implement the early-stopping algorithm presented in Sec.  4.5 and examine empirically how it performs on the real datasets. In Tab. 11, we adopt the following convention: For datasets with $n\leq 300$, we explore $K\in\{2,\dots,30\}$ with tolerance $\epsilon=0.35$. For datasets with $n>300$, we restrict the search to $K\in\{2,\dots,15\}$ with tolerance $\epsilon=0.15$. These parameter choices are not canonical, but rather serve to illustrate different practical regimes. The goal is to demonstrate how Algo. 2 behaves under varying $K$ candidates and tolerance, rather than to prescribe a universally optimal setting.
The results indicate that the proposed upper bound can indeed shorten silhouette-based optimization loops in practice under reasonable tolerance choices. In several cases — notably ceramic, religious-texts, conference-papers, wdbc, ecoli, zoo, iris, and thy — the algorithm halted early while still returning an ASW within the specified tolerance. In other cases, the bound was not tight enough compared to the observed output to trigger early stopping, and the procedure effectively reverted to a full search. This mixed outcome highlights both the promise and the limitations of the approach: when the bound interacts favorably with the empirical ASW, it can reduce computation substantially, but there is no guarantee of savings in every scenario.
Fig. 10 visualizes for the real unlabeled datasets the individual upper bounds in the context of the clusterings obtained by K-means (see Tab. 5.1). It also shows how the upper bound is tightened under size constraints. Fig. 11 and Fig. 12 are constructed in the same way, but for the labeled datasets, without size constraints and with K-medoids instead of K-means. The difference between the two of them is that the latter implements a scan of different $K$ values to maximize the empirical ASW whereas the former uses the ground truth labels to stipulate $K$. Fig. 13 also corresponds to the labeled data and illustrates a more rigorous analysis of the distribution of the individual upper bounds in relation to the $\kappa$ parameter of Algo. 1.
We conclude the discussion by highlighting three key roles of the proposed upper bound.
  1. 1.
    Global upper bound. The unrestricted bound $\mathrm{UB}(\Delta)$ is a ceiling on any achievable ASW for the dataset. It can shorten silhouette-based search loops by enabling early stopping; when tight, it certifies that a globally silhouette-optimal clustering has been found. Even when not tight, it provides essential context for interpreting reported ASW values across algorithms and initializations. Moreover, the magnitude of $\mathrm{UB}(\Delta)$ itself can be seen as a proxy for the clusterability of the dataset: a low upper bound suggests that no clustering can yield well-separated and compact groups, whereas a high upper bound indicates the potential for meaningful structure.
  2. 2.
    Restricted upper bound. Analysts often care about a constrained solution space. By enforcing a minimum cluster size via $\kappa$, the sharper bound $\mathrm{UB}_{\kappa}(\Delta)$ reflects application constraints and yields insights that are more relevant.
  3. 3.
    Individual upper bounds. Our proposed pointwise upper bounds quantify the maximum silhouette each observation can attain. Very low values flag observations that cannot be well integrated under any clustering; such points may merit down-weighting, manual inspection, or exclusion as outliers or noise.
6 Conclusion
In this paper, we propose a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a new tool for analyzing clustering quality. While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a much tighter, data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more meaningful reference value: it often lies well below $1$, reflecting the inherent limitations of the dataset itself.
Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often close-to-sharp and yields valuable insights into the structure and limitations of the data. At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound. For this reason, the bound should be viewed not as a precise target but as an informative ceiling that frames observed ASW scores in a broader context. Even when loose, the bound can shorten optimization loops, offer a first-order estimate of the datasets clusterability, and highlight points that are inherently difficult to cluster.
From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis. We believe this idea opens a promising line of work, where future research should aim at deriving tighter bounds, extending the approach to other clustering validity indices, and exploring additional constraints motivated by application domains. A particularly important direction is to develop theoretical guarantees, for example bounding how far the proposed upper bound can deviate from the true global maximum under certain assumptions on the dataset. This goal could be attained through the derivation of a theoretical and close-to-sharp lower bound on $\mathrm{ASW}(\mathcal{C}^{*},\Delta)$.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
Hugo Sträng: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: visualization, writing, editing, validation.
Acknowledgments
The authors thank Hugo Fernström for his careful proofreading and for offering insightful feedback on a previous version. His constant encouragement greatly contributed to the development of this work.