### v2

# An upper bound of the silhouette validation metric for clustering

Abstract
The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in [1,1][-1,1]. The average silhouette width (ASW) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters. However, the dataset-specific maximum of ASW is typically unknown, and the standard upper limit of 1 is rarely attainable. In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound on the ASW. This bound—often substantially below 1—enhances the interpretability of empirical ASW values by indicating how close a given clustering result is to the best possible outcome on that dataset. It can be used to confirm global optimality, guide the evaluation of clustering solutions, and be refined to incorporate minimum cluster-size constraints for greater practical relevance. Finally, we extend the framework to establish a corresponding bound for the macro-averaged silhouette.
keywords:
Data mining , cluster analysis , internal evaluation metric , silhouette score , average silhouette width , upper bound
[…]
1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [aggarwal2013introduction]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [dinh2025data]. Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [jain1999data].
Among the cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures to assess how well each point fits within its assigned cluster relative to its nearest neighboring cluster [ogsilhouette]. The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$. $\mathrm{ASW}$ values close to 1 signal well‑separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned. Extensive work has examined and refined $\mathrm{ASW}$. [batool2021clustering] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, [john] examined alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets. Intrinsic data characteristics impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$. Comparing $\mathrm{ASW}$ values across datasets with different properties can therefore lead to misguided conclusions about cluster quality.
To address this issue, we derive a data-dependent upper bound on the $\mathrm{ASW}$. This bound supplies essential context: if analysis shows that no partition can exceed, say, an $\mathrm{ASW}$ of 0.30, an empirical result of 0.29 should be regarded as essentially optimal. Such insight prevents wasted effort in pursuit of negligible gains and enables more informed evaluation of clustering quality.
The biggest obstacle in silhouette optimization is the combinatorial explosion of the solution space as the number of samples increases. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion computationally prohibitive. [batool2021clustering] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a data-aware benchmark against which the quality of such locally optimal solutions can be evaluated.
Figure 1: Illustration of the silhouette coefficient calculation.
The main contributions are as follows:
    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures. We show that the bound is nearly tight in many of the cases examined.
    All datasets, preprocessing scripts, bound computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
The rest of this paper is organized as follows. Section 2 reviews related works. Section 3 introduces preliminaries. Section 4 details the proposed method. Section 5 presents the experimental results. Finally, section 6 summarizes the findings and outlines directions for future research.
2 Related work
Silhouette coefficientEnergy & Power Systems[ouyang2024k][rodriguez2024novel][zheng2025practical][duarte2021increasing][eskandarnia2022embedded][zhao2023generalized][koutsandreas2024harnessing][zhou2025recognition][hu2023transformer][alrasheed2024building][diaz2022typologies]Transportation & Mobility[mohd2024passengers][wang2021analyzing][bai2023adaptive][sheng2025multi][guo2025adaptive][passarella2024anomaly][ma2025unravelling][li2025clustering][ansari2021spatiotemporal]Biomedical & Health[jamalpour2025high][sharafoddini2021identifying][ghaderi2023self][gill2024symptom][baier2024nocturnal][kim2022pathological][rustam2022data][jha2021apache][xian2025systematic][bhagat2022dpcf]Food & Agriculture[edris2024identifying][singh2025development][gere2023recommendations]Social & Economic[owoo2025hierarchical][quintero2022analysis][gulma2022new][zhang2025two][aonso2024cluster]Materials & Chemistry[pham2025clustering][bi2025comprehensive][sancho2022cluster][bagirov2023novel][consuegra101unsupervised]Environment & Earth Sciences[nourani2022robust][guo2023combining][yang2022zoning][huang2023dominant][balbi2024hierarchical][biabiany2020design][guan2025novel][gupta2025criteria][zhao2023comparison][deng2025on]Sensor Networks & IoT[kotary2020distributed][kotary2021many][panwar2024distributed][panwar2025distributed][waidyanatha2023regularity]Big Data[qaroush2025maximum][thirumoorthy2021hybrid][song2023novel][chander2023dolphin]Cluster Validity & Selection[khan2025standardizing][khan2025standardization][batool2021clustering][lenssen2024medoid][capo2023fast][gagolewski2021are][bagirov2023finding]Algorithm Development[koren2025preprocessing][silva2025automatic][manochandar2020development][qin2025single][zhang2025two][bai2023adaptive][sheng2025multi][kotary2020distributed][kotary2021many][du2024multi][fu2025gsscl][gan2024discrimination][fu2023gc][enriquez2025performance][gere2023recommendations]Feature & Dim‑Reduction[rustam2022data][dwivedi2025novel][zhou2024ocae][chander2023dolphin]
Figure 2: Taxonomy of how the silhouette coefficient is employed across research fields.
The silhouette coefficient [ogsilhouette] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative. Fig. 1 illustrates the construction. For a point $x_{i}$ in cluster $C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{\ell}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own. Aggregating over points gives the ASW, a dataset-level summary of partition quality.
In practice, ASW is often used for model selection by choosing the number of clusters $K$ that maximizes it. Despite numerous extensions—e.g., to categorical data [dinh2019estimating], fuzzy clustering [lasek2024employing], hierarchical clustering [owoo2025hierarchical], model-based clustering [guan2025novel], spectral clustering [xian2025systematic], spherical clustering [deng2025on], and optimization-based methods [yang2024clustering]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
Other commonly used internal validation indices include the Davies–Bouldin index [dbindex] and the Calinski–Harabasz index [chindex], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
In this work, we focus on the silhouette due to its interpretability and widespread adoption across both research and applied domains. To account for cluster-size imbalance, we additionally employ the macro-averaged silhouette, defined as the unweighted mean of per-cluster silhouette averages. This variant ensures that smaller clusters contribute equally to the overall assessment, providing a more balanced view of clustering quality.
Silhouette-optimization
PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [kaufman90], designed to directly optimize the ASW [pamsil]. Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize ASW. It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in ASW and terminates once no further improvement is possible, ensuring convergence to a local optimum. An efficient implementation of PAMSIL is available in the open-source K-Medoids Python package [Schubert2022]. Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local ASW optima [batool2021clustering].
External validation indices
When ground-truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition. Two of the most commonly used indices are the adjusted Rand index (ARI) [ari] and the adjusted mutual information (AMI) [ami]. Both adjust their underlying similarity measures for chance agreement, ensuring that random labelings yield expected scores near zero.
The ARI extends the classical Rand index by comparing all pairs of samples and counting how consistently they are assigned to the same or different clusters in both partitions. Its normalization accounts for random overlap, producing values in [1,1][-1,1] with $1$ indicating identical partitions. The AMI applies a similar correction to the mutual information (MI) between partitions, normalizing it by the expected MI under random assignments and bounding the result between 0 and $1$.
Both indices are symmetric, easy to interpret, and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
3 Preliminaries
Let $n\in\mathbb{N}$ with $n\geq 4$. We consider a dataset consisting of $n$ objects indexed by the set $\{1,\ldots,n\}$. Each pair of objects is associated with a measure of dissimilarity (or distance), and all pairwise dissimilarities are collected in the dataset’s dissimilarity matrix.
Definition 1 (Dissimilarity matrix).
A matrix $\Delta\in\mathbb{R}^{n\times n}$ is a dissimilarity matrix if
  1. (a)
    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non‑negative),
  2. (b)
    $\Delta_{ii}=0$ for all $i$ ($\Delta$ has zero diagonal), and
  3. (c)
    $\Delta_{ij}=\Delta_{ji}$ for all $i,j$ ($\Delta$ is symmetric).
The triangle inequality is not required. We also exclude matrices containing a row of all zeros.
Figure 3: The first 27 Bell numbers, which enumerate the partitions of an $n$-element set. This rapid growth highlights the combinatorial explosion of clustering possibilities as $n$ increases.
For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order. We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$). In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$. Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
Now, let us formally define the concept of a clustering of the dataset $\{1,\ldots,n\}$.
Definition 2 (Clustering).
Let $K\in\{2,\ldots,n\}$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subsetneq\{i\}_{i=1}^{n}$ for any $I$, and
  1. (a)
    $\bigcup_{I=1}^{K}C_{I}=\{1,\dots,n\}$,
  2. (b)
[…]
  3. (c)
    $C_{I}\neq\varnothing\textnormal{ for all }I\in\{1,\ldots,K\}$.
The total number of clusterings of an $n$-element set grows according to the Bell numbers (Fig. 3), illustrating the combinatorial complexity of the clustering problem.
We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [ogsilhouette]. We employ similar notation.
Definition 3 (Silhouette).
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is 0 if $|C_{I}|=1$; otherwise it is
$s(i|\mathcal{C}_{K},\Delta)\coloneqq\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$
where
$a(i)\coloneqq\frac{1}{|C_{I}|-1}\sum_{j\in C_{I}\setminus\{i\}}\Delta_{ij},\;\;\;\text{and}\;\;\;b(i)\coloneqq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right).$
The average silhouette width (ASW) is defined as
$ASW(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i|\mathcal{C}_{K},\Delta).$
Note that the silhouette width of $i$ may be written as
$s(i|\mathcal{C}_{K},\Delta)=\begin{cases}1-a(i)/b(i)\text{ if }a(i)<b(i),\\ 0\text{ if }a(i)=b(i),\\ b(i)/a(i)-1\text{ if }a(i)>b(i).\end{cases}$
Consequently, $1\leq s(i\,|\,\mathcal{C}_{K},\Delta)\leq 1$.
Individual silhouette widths provide a valuable graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [medoid]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, it is possible that the limitation stems from the intrinsic structure of the dissimilarity data. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$? To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
Definition 4 (Silhouette-optimal clustering).
Let $\Delta$ be the dissimilarity matrix of the dataset $\{1,\ldots,n\}$. We define
$\mathfrak{C}\coloneqq\{\mathcal{C}_{K}:\mathcal{C}_{K}\text{ is a $K$-clustering of $\{1,\ldots,n\}$}\}$
to be the set of all possible clusterings. Any
$\mathcal{C}^{*}\in\operatorname*{argmax}_{\mathcal{C}_{K}\in\mathfrak{C}}\{\mathrm{ASW}(\mathcal{C}_{K},\Delta)\}$
is called a silhouette-optimal clustering.
4 A data-dependent upper bound of the silhouette
We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed. For each data point, we derive a sharp upper bound of its silhouette width. Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and there exists a clustering in which the bound is exactly achieved. Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and the proposed algorithm supports optional minimum cluster size constraints to further tighten the bound.
Consider an arbitrary clustering $\mathcal{C}_{K}$ of the dataset $\{1,\ldots,n\}$ with dissimilarites $\Delta$. We will now formalize the upper bound of the silhouette width for point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)$.
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the remaining data points in the same cluster is not less than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value that is guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning through the $\Lambda$-quotient of point $i$.
Definition 5.
Let $\Lambda\in\mathbb{N}$ with $1\leq\Lambda\leq n-1$. The $\Lambda$-quotient of data point $i$ is
$q(i,\Delta,\Lambda)\coloneqq\begin{cases}\dfrac{n-\Lambda}{\Lambda-1}\cdot\dfrac{\sum_{j=1}^{\Lambda-1}\hat{\Delta}_{ij}}{\sum_{j=\Lambda}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }\Lambda>1,\\[4.30554pt] 1,&\textnormal{if }\Lambda=1.\end{cases}$
Recall the following basic fact.
Lemma 1.
Let $A=\{a_{1},\ldots,a_{N}\}$ be a collection of $N$ real values, and consider a partition of $A$ into subsets $A_{1},\ldots,A_{M}$. Then
$\min_{L}\frac{1}{|A_{L}|}\sum_{a\in A_{L}}a\leq\frac{1}{|A|}\sum_{a\in A}a.$
Proof.
Let $\mu=\frac{1}{|A|}\sum_{a\in A}a$ and $\mu_{L}=\frac{1}{|A_{L}|}\sum_{a\in A_{L}}a$. We have $\mu=\sum_{L=1}^{M}\frac{|A_{L}|}{|A|}\mu_{L}$, and since $\sum|A_{L}|=|A|$, the statement holds. ∎
Proposition 1.
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{i\}_{i=1}^{n}$ with dissimilarites $\Delta$. Let $i$ represent any data point and let $C_{I}\in\mathcal{C}_{K}$ denote the cluster that contains $i$. Furthermore, let
$f(i,\Delta)\coloneqq\min_{1\leq\Lambda\leq n-1}\{q(i,\Delta,\Lambda)\}$
denote a minimal $\Lambda$-quotient. Then
$s(i|\mathcal{C}_{K},\Delta)\leq 1-f(i,\Delta).$
Proof.
Suppose $q(i,\Delta,2)>1$. Then $(n-2)\hat{\Delta}_{i1}>\sum_{j=2}^{n-1}\hat{\Delta}_{ij}$, which contradicts $\hat{\Delta}_{i1}\leq\hat{\Delta}_{ij}$ for $j\geq 2$. Hence, $0\leq f(i,\Delta)\leq 1$. This proves the statement if $|C_{I}|=1$ and for cases $a(i)\geq b(i)$ when $|C_{I}|>1$. Consider $q(i,\Delta,|C_{I}|)$. It is clear that $\sum_{j=1}^{|C_{I}|-1}\hat{\Delta}_{ij}/(|C_{I}|-1)\leq a(i)$. It remains to argue that $\sum_{j=|C_{I}|}^{n-1}\hat{\Delta}_{ij}/(n-|C_{I}|)\geq b(i)$. By Lemma 1, we have
$\frac{1}{n-|C_{I}|}\sum_{j=|C_{I}|}^{n-1}\hat{\Delta}_{ij}$ $\geq\frac{1}{n-|C_{I}|}\sum_{j\in\bigcup_{J\neq I}C_{J}}\Delta_{ij}$
$\geq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right)=b(i).$
Thus, $f(i,\Delta)\leq q(i,\Delta,|C_{I}|)\leq a(i)/b(i)$ and $1-f(i,\Delta)\geq 1-a(i)/b(i)$. ∎
Taking the mean of all silhouette width upper bounds from Proposition 1 results in the following inequality, which is the main observation of this paper:
$ASW(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
We remark that the right-hand side in 1 depends only on the underlying data $\Delta$. Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$. Moreover, since any $\Lambda$-quotient naturally corresponds to a $2$-clustering of the dataset, the bound in Proposition 1 is indeed sharp.
Finally, imposing the narrower definition
$f_{\kappa}(i,\Delta)\coloneqq\min_{\kappa\leq\Lambda\leq n-\kappa}\{q(i,\Delta,\Lambda)\}$
for some $\kappa\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$. Often, as we shall see, this leads to strictly tighter limits.
4.1 Algorithm
Input: $\Delta\in\mathbb{R}^{n\times n}$;
// Dissimilarity matrix
1
Input: $\kappa$;
// Smallest allowed cluster size, $1\leq\kappa\leq\lfloor n/2\rfloor$
2
Output: $A\in\mathbb{R}^{n}$;
// Array of individual upper bounds
3
4 Initialize array $A\leftarrow[\;]$ ;
5 Compute $\hat{\Delta}$ from $\Delta$;
6 for $i=1,\ldots,n$ do
7 $y\leftarrow\sum_{j=\kappa}^{n-1}\hat{\Delta}_{i,j}$ ;
8 $x\leftarrow 0$ ;
 $q\leftarrow 1$ ;
 // Initialize $\Lambda$-quotient
9 if $\kappa>1$ then
10    $x\leftarrow\sum_{j=1}^{\kappa-1}\hat{\Delta}_{i,j}$ ;
11    $q\leftarrow\frac{x}{\kappa-1}\big/\frac{y}{n-\kappa}$;
12
13 for $\Lambda=\kappa+1,\ldots,n-\kappa$ do
14    $x\leftarrow x+\hat{\Delta}_{i,\Lambda-1}$;
15    $y\leftarrow y-\hat{\Delta}_{i,\Lambda-1}$;
16    $q_{\text{candidate}}\leftarrow\dfrac{x}{\Lambda-1}\big/\dfrac{y}{n-\Lambda}$;
17    if $q_{\text{candidate}}<q$ then
       $q\leftarrow q_{\text{cand}}$ ;
       // Update minimal $\Lambda$-quotient
18
19
20  Append 1q1-q to AA;
21
return AA
Algorithm 1 Silhouette upper bound (pointwise)
The workflow for computing the proposed upper bound is illustrated in Fig.4. Following this workflow, Algo. 1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the solution space $\min_{I}|C_{I}|\geq\kappa$. The array $A$ collects these values across all points, and its average equals the dataset-level upper bound.
Time complexity
The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\mathrm{ASW}$, which takes $\mathcal{O}(n^{2})$ time [medoid].
Output summaries
Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo. 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). We refer to this as the global upper bound. Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects. Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space. We will refer to $\texttt{UB}_{\kappa}(\Delta)$ as an adjusted upper bound.
Remark
The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$. Its tightness is data-dependent.
Use cases
In theory, when sharp, the upper bound $\texttt{UB}(\Delta)$ can be used to confirm that an empirical ASW has reached the global optimum. Even when loose, it increases the interpretability of empirical measurements.
In real-world scenarios where the global upper bound appears loose, we speculate that adjusted versions can provide more practical use.
The individual silhouette width upper bounds generated by Algo. 1 provide a neat complement to standard silhouette visualizations, where the point-wise silhouette widths are typically displayed next to each other in a bar graph.
In the Experiments section, we examine these use cases.
Figure 4: Workflow for computing the proposed $\mathrm{ASW}$ upper bound.
4.2 Extension to the macro-averaged silhouette
The standard $\mathrm{ASW}$ can produce misleading results when the clusters vary greatly in size. In [john], John Pavlopoulos, Georgios Vardakas, and Aristidis Likas show that the so-called macro-averaged silhouette is more robust to cluster-size imbalance. For clustering $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$, the macro-averaged silhouette is defined as
$S_{macro}(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{K}\sum_{I=1}^{K}\frac{1}{|C_{I}|}\sum_{i\in C_{I}}s(i).$
In practical scenarios where clusters are imbalanced and small groups also matter, this aggregation strategy is more reasonable. In this subsection, we use Proposition 1 to construct an upper bound for this silhouette variant in a constrained solution space.
Consider an empirical clustering result $\widehat{\mathcal{C}}_{K}=\{\widehat{C}_{I}\}_{I=1}^{K}$. We narrow our scope in this context to include all clusterings whose cluster sizes match those of $\widehat{\mathcal{C}}_{K}$. This narrower solution space is still generally astronomically large. Thus, an efficiently computable data-dependent upper bound is of interest as a relevant benchmark. Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$. Moreover, let $\{\texttt{ub}_{\sigma(i)}\}_{i=1}^{n}$ denote the upper bounds sorted in ascending order, and let $\{N_{I}\}_{I=1}^{K}$ denote the assumed cluster sizes in descending order. With $N_{0}\coloneqq 1$, it follows from the rearrangement inequality that
$S_{macro}(\mathcal{C}_{K},\Delta)$ $\leq\frac{1}{K}\sum_{I=1}^{K}\sum_{i\in C_{I}}\frac{1}{|C_{I}|}\texttt{ub}_{i}$
$\leq\frac{1}{K}\sum_{I=1}^{K}\frac{1}{N_{I}}\sum_{i=\sum_{J=0}^{I-1}N_{J}}^{\sum_{J=1}^{I}N_{J}}\texttt{ub}_{\sigma(i)}.$
The last expression is independent of any particular cluster assignment and therefore provides a valid upper bound on the macro-averaged silhouette across all clusterings within the constrained solution space. This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated. Even in the absence of such assumptions, the resulting bound serves as a sharper and more informative benchmark for evaluating empirical clustering results.
5 Experiments, results & discussion
Table 1: Real-world datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
Dataset Title Instances Features Classes Field Metric
Ceramic Chemical Composition of Ceramic Samples 88 17 2 Physics Manhattan
Customers Wholesale customers 378 7 2 Business Cosine
Dermatology Dermatology 366 34 6 Health and Medicine Cosine
Heart Statlog Statlog (Heart) 231 13 2 Health and Medicine Euclidean
Optdigits Optical Recognition of Handwritten Digits 2296 64 10 Computer Science Euclidean
Rna Gene expression cancer RNA-Seq 808 20531 5 Biology Euclidean
Wdbc Wisconsin Diagnostic Breast Cancer 382 30 2 Health and Medicine Euclidean
Wine Wine 178 13 3 Chemistry Manhattan
Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound.
Dataset $K$ ASW $\texttt{UB}(\Delta)$ Worst-case rel. error
400-64-5-6 5 $0.249$ $0.376$ $0.339$
400-64-2-2 2 $0.673$ $0.673$ $0.000$
400-128-7-3 7 $0.522$ $0.566$ $0.078$
1000-300-5-2 5 $0.668$ $0.678$ $0.015$
10000-32-20-2 20 $0.626$ $0.774$ $0.192$
10000-1024-20-4 20 $0.417$ $0.454$ $0.081$
We evaluate the ASW upper bound by applying it to a variety of datasets, both synthetic and real-world. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
An essential part of the evaluation is to establish a nearly tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$. To this end, we use PAMSIL clustering to obtain an empirical $\mathrm{ASW}$. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies. PAMSIL is a suitable algorithm since its objective is to maximize ASW.
Given an empirical ASW value, say $\widehat{S}$, the performance metric of interest is the relative error, defined as
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)},$
where $C^{*}$ denotes a clustering that achieves the global maximum. Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical ASW is guaranteed to be within $\epsilon$ of the true maximum.
Synthetic datasets
We follow [gaussian] in notation and approach. Synthetic datasets are generated using the make_blobs function from scikit-learn [scikit]. Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion. Finally, for these datasets, we use Euclidean distance.
Real-world datasets
We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [ucidata], commonly used in clustering and classification research. Table 1 summarizes their key characteristics along with the distance metrics employed. Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate. Each dataset provides a known number of classes, which we use to determine the number of clusters $K$. When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures. In that way, we ensure that our experiments are conducted in natural clustering contexts.
Table 3: Experimental results on real data. Comparison of clustering performance and the proposed upper bound.
Performance of PAMSIL clustering
Comparison with
global upper bound
Comparison with
adjusted upper bound
Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI ASW $\texttt{UB}(\Delta)$ Worst-case rel. error $\texttt{UB}_{\kappa}(\Delta)$ Worst-case rel. error
Ceramic 2 44 $1.000$ $1.000$ $0.358$ $0.621$ $0.423$ $0.403$ $0.112$
Customers 2 170 $0.394$ $0.366$ $0.400$ $0.951$ $0.579$ $0.574$ $0.303$
Dermatology 6 21 $0.850$ $0.875$ $0.455$ $0.836$ $0.456$ $0.704$ $0.354$
Heart Statlog 2 83 $0.447$ $0.349$ $0.161$ $0.600$ $0.731$ $0.279$ $0.422$
Optdigits 10 50 $0.803$ $0.819$ $0.180$ $0.620$ $0.709$ $0.463$ $0.611$
Rna 5 77 $0.225$ $0.419$ $0.464$ $0.323$ $0.303$
Wdbc 2 46 $0.427$ $0.386$ $0.348$ $0.621$ $0.440$ $0.455$ $0.236$
Wine 3 53 $0.884$ $0.864$ $0.315$ $0.649$ $0.514$ $0.450$ $0.298$
Figure 5: Silhouette plots for the eight real-world datasets. PAMSIL clustering with $K$ equal to the ground truth number of classes; ASW compared against the global upper $\texttt{UB}(\Delta)$ and the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
Figure 6: PAMSIL performance for varying $K$ and comparison with upper bounds.
Results
In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering. For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score. For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL ASW is within 8 $\%$ of the global maximum, that is, essentially optimal. For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical ASW of $0.249$. In this case, it remains unclear whether further substantial improvements of the ASW are possible. However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1].
Table 3 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds. In these cases, the global upper bounds provide information that is less actionable. It is unclear whether the reason for this is that the attained ASW values are far from $\mathrm{ASW}(C^{*},\Delta)$, or that the upper bounds are somewhat loose, or both. However, the adjusted upper bounds $\texttt{UB}_{\kappa}(\Delta)$ are considerably tighter for most datasets. Here, $\kappa$ is set to the minimum cluster size in the PAMSIL solution. For datasets Ceramic, Customers, Rna, Wdbc and Wine, $\texttt{UB}_{\kappa}(\Delta)$ proves that the ASW-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums.
Fig. 5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds. For the adjusted upper bounds, $\kappa$ is set to the smallest cluster size in the generated solution. This method provides a visual comparison of the empirical silhouette and the highest possible value at the data point level.
In Fig. 6 we show, for each real-world dataset and for each $K\in\{2,\ldots,15\}$, the attained ASW value compared to the global ASW upper bound. Additionally, at each $K$, we compare the ASW to the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$ with $\kappa$ set to the minimum cluster size in the corresponding solution. Lastly, we include, at each $K$, the recorded macro-averaged silhouette and compare this value with the upper bound derived in Section 4.2.
Discussion
We summarize with some reflections on the practical relevance of the proposed upper bound.
Replacing the standard $\mathrm{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\mathrm{ASW}$ values. However, when the empirical $\mathrm{ASW}$ lies far below the upper bound, it remains unclear whether further improvement is actually possible. We suspect that this gap often persists for the global upper bound, even when the empirical $\mathrm{ASW}$ is close to $\mathrm{ASW}(C^{*},\Delta)$. In many real-world applications, however, analysts impose minimum cluster-size constraints to avoid overly small or degenerate clusters. By incorporating such constraints through $\kappa$, the refined bound $\texttt{UB}_{\kappa}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice. As demonstrated in Table 3, $\texttt{UB}_{\kappa}(\Delta)$ ensures that the PAMSIL solutions are within 30 $\%$ of the optimum for five of the examined datasets, indicating strong empirical performance under realistic conditions. In such cases, the bound can be used to conclude that further exploration would, at best, yield only marginally better results.
6 Conclusion
In this paper, we proposed a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a complementary tool for analyzing clustering quality. While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more meaningful reference value. It often lies well below $1$, reflecting the inherent limitations of the dataset itself.
Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often nearly tight and yields valuable insights into the structure and limitations of the data. At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound. For this reason, the bound should not be viewed as a precise target but as an informative ceiling that frames observed ASW scores in a sharper context.
From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis. We believe this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices. An important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its tightness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters. Such insights would clarify in which contexts the bound provides the greatest practical value.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
Hugo Sträng: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: visualization, writing, editing, validation.
Acknowledgments
The authors would like to thank Hugo Fernström for his careful proofreading and thoughtful feedback on an earlier version of this manuscript. His continuous encouragement has been greatly appreciated throughout the development of this work.
We are also grateful to several researchers for their valuable comments and suggestions on a previous version of the paper.